When the air resistance is ignored, the horizontal range R of a projectile is given by R(8)= 9 sin 20 where is the constant initial velocity, g is the acceleration due to gravity, and is the angle of elevation or departure. Find the maximum range of projectile.


1
Expert's answer
2021-11-15T17:55:36-0500

In general, the horizontal range "R" of a projectile is given by "R(\alpha)=\frac{v_0^2\cdot\sin2\alpha}{g}" .


Tthe maximum range of projectile will be when "\sin2\alpha=1\to \alpha=45°"


"R_{max}(\alpha)=R_{max}(45°)=\frac{v_0^2}{g}" .